OpenStudy (anonymous):

Help please, will give medal :D A ramp leads up to a building. The top of the ramp is 6 feet above the ground, and the bottom is 16 feet from the building as shown. What is the length of the ramp? Round your answer to the nearest tenth of a foot.

OpenStudy (anonymous):



OpenStudy (anonymous):

\[h^2=16^2+6^2\\ h=\sqrt{16^2+6^2}\] whatever that is